OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x + 3y = 15?

OpenStudy (anonymous):

parallel lines have equal slopes !!

OpenStudy (anonymous):

3y = 15-5x, y = 5-(5/3)x so m = (5/3)